the product of three consecutive integers is 120. what is their sum?
Found 2 solutions by richard1234, rashedbiman:
Answer by richard1234(7193)

990 = 9*10*11, sum is 30.

Answer by rashedbiman(2)

Guys ths is very easy solution: good luck
ans.
120/2=60
again, 60/2=30
again, 30/2=15
again, 15/3=5
so here all the produnts is = 2*2*2*3*5
= 4*6*5
so the consecutive integers= 4*5*6
sum= 4+5+6
= 15
goodluck